# MARFT: Multi-Agent Reinforcement Fine-Tuning

### v3

@@ 4.2.1 Base Algorithm
The detailed algorithm implementation is illustrated as algorithm 3, and the details are demonstrated in the Appendix A with highlighted comments as explanations. For the design of the value function $V_{\phi}$, we base it with a frozen transformer module, which proceeds the sequence information and outputs the final hidden state vector of the sequence, and add an extra trainable multilayer perceptron (MLP) with parameters $\phi$ to map the hidden vector to a specific value. A common instantiation of the transformer module can be the base model of an agent. For the design of the agents, we usually resort to LoRA fine-tuning to avoid overwhelming pressure on computing resources. If the agents share the same base model, we can assign separate but same-structured LoRA adapters to them, but if the agents have different base models, even different modalities, we then have to give them distinguished LoRA adapters.
@@
6                   format roll-out observation $\hat{o}^{i}_{t}$ by concatenating agent $i$’s profile $\text{inst}^{i}$ and environment observation $o_{t}^{i}$
7                   generate action $a^{i}\sim p_{\pi^{i}_{\theta^{i}}}(a^{i}|\hat{o}^{i},\bm{a}^{\mathcal{N}(i)}),\;\mathcal{N}(i)=\left\{j\neq i\mid\mathcal{D}(a^{j},a^{i})=1\right\}$
@@
9            $\bm{o}_{t+1}\sim\mathcal{T}(\cdot|\bm{o}_{t},\bm{a}_{t})$
10             $r_{t}=R(\bm{o}_{t},\bm{a}_{t})$
11             $\mathcal{D}\leftarrow\mathcal{D}\cup\{\left(\hat{\bm{o}}_{t},\bm{a}_{t},r_{t},\bm{o}_{t+1}\right)\}$
12
@@ 11             $\mathcal{D}\leftarrow\mathcal{D}\cup\{\left(\hat{\bm{o}}_{t},\bm{a}_{t},r_{t},\bm{o}_{t+1}\right)\}$
16             sample a batch =$\mathcal{B}=\{(\hat{\bm{o}}_{t},\bm{a}_{t},r_{t},\bm{o}_{t+1},\hat{A},\hat{V})\}\sim\mathcal{D}$
17             update $\phi$ by minimizing $\sum_{n=1}^{N}\left\|V_{\phi}(\hat{\bm{o}}_{n})-\hat{V}_{n}\right\|^{2}$
@@
Algorithm 3 Multi-Agent Reinforcement Fine-Tuning (MARFT)
@@
${\text{Norm}}\left(\text{log}P_{\pi^{i}_{\theta^{i}}}\left(a^{i}|\hat{o}^{i},\bm{a}^{1:i-1}\right)\right)=\frac{\text{log}P_{\pi^{i}_{\theta^{i}}}\left(a^{i}|\hat{o}^{i},\bm{a}^{i-1}\right)}{Z}.$ (26)
@@ 4.2.2 Token-level Adaptations
For the implementation of token-level MARFT, the overall procedure mirrors that of Algorithm 3, with the primary modification being the storage of all token logits and values, rather than just the joint action probabilities and values, during the trajectory rollout phase. Subsequently, the token-level Bellman backup is utilized to determine the target values for the critic network and to calculate the TD errors. Additionally, GAE (Schulman et al., 2018) is employed to estimate the advantage values for each token. The optimization objectives remain consistent with those outlined in the general framework.
@@ 5 Experiments
To verify the feasibility and effectiveness of the MARFT paradigm, we implement specific and naive instantiations [^4] MARFT itself is a new paradigm, not a specific implementation. In this section, we refer to action-level and token-level MARFT as specific naive implementations based on PPO without other dedicated techniques. that follows its general principles and conduct a series of experiments [^5] All experiments are conducted on machines equipped with NVIDIA H100 GPUs (80GB), using PyTorch 2.4 and CUDA 12.1. Each experiment utilizes only 1 GPU., including main experiments of tuning LaMAS in math problem-solving environments, and ablation experiments on optimization granularity, agent population, and value pertaining. We demonstrate the learning dynamics for further analysis. More experiments and experiment details are presented in Appendix B.
@@ 5.1 Setups
Task Environment
We constructed a task environment for MATH problem-solving. The environment is supported by specific datasets and reset by randomly sampling a (problem, answer) pair from the dataset. The LaMAS takes actions, and the environment verifier checks the correctness. The reward is a coarse-grained binary signal: 1 for correct answers and 0 otherwise.
@@
We use MATH (Hendrycks et al., 2021) and CMATH (Wei et al., 2023) for English and Chinese math problem-solving, respectively. The environments are built using the MATH train set (7.5k entries, 7 types), the CMATH train set (600 entries), and a verifier. Evaluation is conducted in-domain and cross-domain using MATH500 (500 entries) and the CMATH (1098 entries) test set, and out-of-domain using the GSM8K (Cobbe et al., 2021) test set (1319 entries).
@@
We use Qwen2.5-Coder-3B-Instruct (Yang et al., 2024a; Hui et al., 2024) as the base model. Solo represents a single-agent setup, while Duo represents a LaMAS with two agents (Reasoner $\rightarrow$ Actor) collaborating on problem-solving. Each agent is equipped with a dedicated LoRA adapter. Profile configurations for these experiments are detailed in Appendix B.3.
@@ 5.2 Learning Dynamics
Figure 4 shows the training dynamics of MARFT-tuned Duo in environments supported by MATH and CMATH. The episodic return is the correctness of solving the math problem. The average step reward is calculated by dividing the episodic return by the steps taken to solve the problem.
(a) Episodic Return (MATH).
(b) Average Step Reward (MATH).
(c) Value Loss (MATH).
(d) Episodic Return (CMATH).
(e) Average Step Reward (CMATH).
(f) Value Loss (CMATH).
Figure 4: Learning dynamics during Duo MARFT (computed over 3 random seeds).
We observe that in both the MATH and CMATH settings, the episodic return exhibits a clear improvement, up to about 8% over the baseline performance. However, during the early stages of training, the episodic return demonstrates noticeable oscillations. This instability can be explained by the behavior of the value loss curves, which display high losses during the initial training phase. Besides, we observe a constant improvement in average step reward, which means Duo LaMAS not only tries to achieve high episodic return but also learns to solve the problem more efficiently.
@@ 5.3 Evaluation Results
On top of the learning dynamics of MARFT in the environment supported by the train set, we also conduct evaluations on the test set. Table 3 presents the performance of Solo, Duo, and MARFT-tuned Duo on the MATH500, CMATH, and GSM8K test sets, including in-domain and out-of-domain assessments.
Solo vs. Duo
Comparing Solo and Duo, we observe that Duo significantly outperforms Solo on the MATH500 benchmark by approximately 7%. However, their performance is comparable on CMATH and GSM8K, likely because these datasets consist of relatively simpler problems. In such cases, a more complex solving process may hinder performance. These results suggest that a LaMAS like Duo is more effective than a single agent in solving complex mathematical problems.
Table 3: Evaluation results of Solo, Duo and MARFT-tuned Duo. Duo-MATH and Duo-CMATH mean MARFT-tuned Duo in the environments supported by MATH train set and CMATH train set.
Benchmark Solo Duo Duo-MATH Duo-CMATH
MATH500 40.4% 47.2% 49.8% 49.4%
CMATH 81.4% 81.1% 83.0% 83.2%
GSM8K 76.8% 77.3% 78.7% 79.4%
Duo vs. MARFT-Tuned Duo
On in-domain benchmarks, MARFT enhances Duo’s performance by around 2% to 3%. Additionally, MARFT improves Duo’s effectiveness in out-of-domain settings, demonstrating its capability to foster collaborative behavior within Duo. This indicates that MARFT can further augment the performance of multi-agent systems.
Out-of-Domain Evaluation
Duo-MATH and Duo-CMATH, fine-tuned on English and Chinese datasets respectively, both generalize well across domains: Duo-CMATH scores 49.4% on MATH500, and Duo-MATH reaches 83.0% on CMATH. Both also improve performance on GSM8K, indicating that MARFT enhances not only in-domain accuracy but also cross-lingual and out-of-domain generalization.
@@ 5.4.1 Optimization Granularity
Figure 5 (Figure 7 in Appendix) shows Duo MARFT with different optimization granularity, i.e., action-level and token-level. From both subfigures, we see that action-level MARFT is overall better than token-level MARFT. Specifically, action-level MARFT achieves a higher episodic return than token-level MARFT in the end and has a more stable and faster improvement. [^6] We early stop some experiments, since as the training goes on, LaMAS may explore invalid tokens, leading to a burst of entropy and collapse of training. Currently, this is a common and unavoidable phenomenon in RL for LLMs, though it can be alleviated by using some tricks (Yu et al., 2025; Yue et al., 2025). This kind of superiority of action-level over naive token-level can be theoretically deduced by the inconsistency of their optimality, which is caused by the different MDP (Wen et al., 2024b).
@@
5.4.2 Agent Number (Solo vs. Duo)
Table 4: The performance ceilings of Solo and Duo during MARFT.
Environment Solo Duo
MATH $\sim$ 45% $\sim$ 53%
CMATH $\sim$ 85% $\sim$ 87%
We also conduct MARFT with the number of agents set to 1, which reduces to single-agent PPO. The asymptotically best performances of both Solo and Duo are shown in Table 4 (Figure 8 in the Appendix), from which we observe that Solo’s peak performance reaches approximately 45% on MATH and 85% on CMATH—about 8% and 2% lower, respectively, than Duo’s during training, and the harder the problem sets are, the more obvious the differences are. This result suggests that LaMAS possesses a higher performance ceiling than a single agent, even in this relatively constrained mathematical setting, as opposed to more complex agentic scenarios.
@@ 5.5 Future Experiments
Given the preliminary nature of our experiment on the only MATH (Hendrycks et al., 2021) and CMATH (Wei et al., 2023), we outline a series of future experiments to further explore the capabilities and robustness of MARFT:
  • MARFT on coding tasks: Introduce MARFT to handle coding tasks and benchmarks such as LiveCodeBench (Jain et al., 2024) and CodeForces [^7] https://codeforces.com/. This setup will test MARFT’s ability to manage iterative problem-solving and reflection in complex coding environments.
@@ 7 Open Problems
Developing engineering solutions that integrate LLMs and MARL is highly challenging. While there exist user-friendly, efficient, and scalable frameworks such as OpenRLHF (Hu et al., 2024) for LLM fine-tuning and MALib (Zhou et al., 2023) for MARL training, a comprehensive framework specifically designed for MARFT is notably absent. The creation of such a framework would require significantly more effort than developing either of the aforementioned frameworks individually. This is because it must effectively integrate both LLMs and MARL components, each of which presents its own set of complexities. While this article introduces a MARFT framework, it is still in its early stages and offers significant room for further development and refinement.
@@ 8 Conclusion
In this article, we propose a novel paradigm termed Multi-Agent Reinforcement Fine-Tuning (MARFT), together with a brand-new POMDP called Flex-POMDP to better align with the LaMAS optimization in realistic scenarios, and present a universal algorithmic framework, MARFT, that integrates RFT, LaMAS, and MARL to advance the capabilities of LaMAS systems. MARFT addresses several fundamental challenges in adapting conventional MARL to LaMAS, including asynchronous agent interactions, profile-aware architecture design, heterogeneity across agents, and the dynamic organization of multi-agent systems. Through the fusion of RFT with sequential decision-making paradigms, trust-region optimization techniques, and token-level adaptation mechanisms, MARFT enables efficient coordination and policy refinement while preserving the foundational capabilities of pretrained LLMs.
@@ Appendix A Detailed Framework Code Explanation
Abstract action-level MARFT training code in the trainer. (For token-level training, we just need to modify the policy loss computation to take token logits as input):
@@
37 critic_grad_norm = nn.utils.clip_grad_norm_(self.mas.critic.parameters(), self.max_grad_norm)
@@
70 policy_grad_norm = nn.utils.clip_grad_norm_(self.mas.agents.parameters(), self.max_grad_norm)
@@
76 policy_grad_norm = nn.utils.clip_grad_norm_(self.mas.agents.parameters(), self.max_grad_norm)
@@ B.2 Hyperparameter Configurations for Experiments
Table 6: Hyperparameter configuration for token-level MARFT in Figure 5.
@@
Table 7: Hyperparameter configuration for tuning Solo in Table 4.
Hyperparameter Value Hyperparameter Value Hyperparameter Value
MAS learning rate 3e-6 critic learning rate 1e-5 use gae True
rollout threads 2 num mini-batch 1 ppo epoch 1
max horizon 2 hidden size 64 episode length 8
optim eps 1e-5 critic loss coef 1 entropy coef 1e-3
gamma 0.99 gae gamma 0.95 max grad norm 0.5
context window 1536 max new tokens 512 clip param 0.2

### v4

@@ 4.2.1 Base Algorithm
The detailed algorithm implementation of MARFT-A is illustrated as algorithm 3, and the details are demonstrated in the Appendix A with highlighted comments as explanations. For the design of the value function $V_{\phi}$, we base it with a frozen transformer module, which proceeds the sequence information and outputs the final hidden state vector of the sequence, and add an extra trainable multilayer perceptron (MLP) with parameters $\phi$ to map the hidden vector to a specific value. A common instantiation of the transformer module can be the base model of an agent. For the design of the agents, we usually resort to LoRA fine-tuning to avoid overwhelming pressure on computing resources. If the agents share the same base model, we can assign separate but same-structured LoRA adapters to them, but if the agents have different base models, even different modalities, we then have to give them distinguished LoRA adapters.
@@
6         generate action $a^{i}\sim p_{\pi^{i}_{\theta^{i}}}(a^{i}|s,\bm{a}^{1:i-1})$
@@
8    $s_{t+1}\sim\mathcal{T}(\cdot|s_{t},\bm{a}_{t})$
9    $r_{t}=R(s_{t},\bm{a}_{t})$
10    $\mathcal{D}\leftarrow\mathcal{D}\cup\{\left(s_{t},\bm{a}_{t},r_{t},s_{t+1}\right)\}$
11
@@ 10    $\mathcal{D}\leftarrow\mathcal{D}\cup\{\left(s_{t},\bm{a}_{t},r_{t},s_{t+1}\right)\}$
15      sample a batch $\mathcal{B}=\{(s_{t},\bm{a}_{t},r_{t},s_{t+1},\hat{A},\hat{V})\}\sim\mathcal{D}$
16      update $\phi$ by minimizing $\sum_{n=1}^{N}\left\|V_{\phi}(s_{n})-\hat{V}_{n}\right\|^{2}$
@@
Algorithm 3 Action-level Multi-Agent Reinforcement Fine-Tuning (MARFT-A)
@@
${\text{Norm}}\left(\text{log}P_{\pi^{i}_{\theta^{i}}}\left(a^{i}|s,\bm{a}^{1:i-1}\right)\right)=\frac{\text{log}P_{\pi^{i}_{\theta^{i}}}\left(a^{i}|s,\bm{a}^{i-1}\right)}{Z}.$ (22)
@@ 4.2.2 Token-level Adaptations
For the implementation of MARFT-T, the overall procedure mirrors that of Algorithm 3, with the primary modification being the storage of all token logits and values, rather than just the joint action probabilities and values, during the trajectory rollout phase. Subsequently, the token-level Bellman backup is utilized to determine the target values for the critic network and to calculate the TD errors. Additionally, GAE (Schulman et al., 2018) is employed to estimate the advantage values for each token. The optimization objectives remain consistent with those outlined in the general framework.
@@ 5 Experiments
To verify the feasibility and effectiveness of the MARFT paradigm, we implement MARFT-A and MARFT-T that follows general principles of MARFT and conduct a series of experiments, including main experiments of fine-tuning LaMAS in math problem-solving and coding environments. We demonstrate the learning dynamics and evaluation results [^5] The mean and standard error of all curves are computed over 3 random seeds. The mean and standard error of evaluation results are calculated over 10 random seeds with temperature 0.1. for further analysis. More experiments and experiment details are presented in Appendix B.
@@ 5.1 Setups
Task Environments
The environments are supported by specific datasets and reset by randomly sampling a (problem, answer, or test cases) pair from the dataset. The LaMAS takes actions, and the environment verifier checks the correctness. The reward in math problem-solving environment is 1 for correct answers and 0 otherwise, and the reward in coding environment is the test case pass rate.
@@
For math problem-solving, we train on the MATH (Hendrycks et al., 2021) (7.5k entries) and CMATH (Wei et al., 2023) (600 entries) training sets. We evaluate on MATH500, the CMATH test set, and the out-of-domain GSM8K (Cobbe et al., 2021) test set. For coding, we use the CodeForces (Penedo et al., 2025) dataset with 1339 training and 377 test entries.
@@
For math experiments, we use Qwen2.5-Coder-3B-Instruct (Yang et al., 2024a; Hui et al., 2024) as the base model [^6] The models were specifically chosen to avoid data leakage.. Solo represents a single-agent setup, while Duo represents a LaMAS with two agents (Reasoner $\rightarrow$ Actor) collaborating on problem-solving. For coding experiments, we use Qwen2.5-3B-Instruct (Yang et al., 2024a; Hui et al., 2024) as the base model. On top of the Duo setting used in the math scenario, we add a Trio that represents a LaMAS with three agents (Reasoner $\rightarrow$ Coder $\rightarrow$ Reviewer). Each agent is equipped with a dedicated LoRA adapter. Profile configurations for these experiments are detailed in Appendix B.3.
@@ 5.2 Learning Dynamics
Figure 4 shows the training dynamics of MARFT-A of Duo in environments supported by MATH and CMATH. The episodic return (ER) is the correctness of solving the math problem. The average step reward (ASR) is calculated by dividing the episodic return by the steps taken to solve the problem.
(a) ER (MATH).
(b) ASR (MATH).
(c) ER (CMATH).
(d) ASR (CMATH).
Figure 4: Learning dynamics of MARFT-A of Duo on MATH and CMATH.
(a) Duo.
(b) Trio.
Figure 5: Learning dynamics of MARFT-A on CodeForces.
We observe that in math problem-solving environments, the episodic return exhibits a clear improvement up to about 8 p.p.777p.p. is the abbreviation for percentage points. (~18.45%). However, during the early stages of training, the episodic return demonstrates noticeable oscillations. This instability can be explained by the behavior of the value loss, which display high losses during the initial phase. Besides, we observe a constant improvement in average step reward, which means Duo not only tries to achieve high episodic return but also learns to solve the problem more efficiently.
Figure 5 shows the training dynamics of MARFT-A of Duo and Trio in coding environments built by CodeForces. We observe that the scores undergo a certain oscillation in the beginning and then increase stably from about 220 to about 260 (~18.18%) and 280 (~27.27%) respectively for Duo and Trio.
@@ 5.3 Evaluation Results
On top of the learning dynamics of MARFT-A in the environment supported by the train set, we also conduct evaluations on the test set. Table 3 presents the performance of Solo, Duo, and MARFT-A-tuned Duo on the MATH500, CMATH, and GSM8K test sets, including in-domain and out-of-domain assessments. And Table 4 gives a comparison of total scores between vanilla LaMAS and the MARFT-A-tuned ones on the test set.
Solo vs. Multi-Agent
For mathematical tasks, by comparing Solo and Duo, we observe that Duo significantly outperforms Solo on the MATH500 benchmark by approximately 7.5 p.p. (~20.58%). However, their performance is comparable on CMATH and GSM8K, likely because these datasets consist of relatively simpler problems. In such cases, a more complex solving process may hinder performance. These results suggest that a LaMAS like Duo is more effective than a single agent in solving complex mathematical problems. For coding tasks, Solo and Duo perform close to each other, but Trio falls behind by about one point, which proves that overcomplicating a system for a task with modest difficulty can harm the system’s performance.
Table 3: Evaluation results (in percentages) of Solo, Duo and MARFT-A-tuned Duo on mathematical benchmarks. MARFT-A Duo-MATH and MARFT-A Duo-CMATH mean MARFT-A-tuned Duo in the environments supported by MATH train set and CMATH train set.
Benchmarks Solo Duo MARFT-A
Duo-MATH Duo-CMATH
MATH500 $36.44\pm 1.35$ $43.9\pm 1.09$ $47.14\pm 1.31$ $47.18\pm 0.10$
CMATH $80.24\pm 0.62$ $79.00\pm 0.98$ $81.26\pm 0.55$ $81.82\pm 0.85$
GSM8K $68.45\pm 1.68$ $4.23\pm 0.63 77$$7.24\pm 0.94$ $76.76\pm 0.55$
Vanilla vs. MARFT-A
For mathematical tasks, on in-domain benchmarks, MARFT-A enhances Duo’s performance by about 3 p.p.. Additionally, MARFT-A improves Duo’s effectiveness in out-of-domain settings, demonstrating its capability to foster collaborative behavior within Duo. This indicates that MARFT-A can further augment the performance of multi-agent systems. Duo-MATH and Duo-CMATH, fine-tuned on English and Chinese datasets respectively, both generalize well across domains: Duo-CMATH scores 49.4% on MATH500, and Duo-MATH reaches 83.0% on CMATH. Both also improve performance on GSM8K, indicating that MARFT enhances not only in-domain accuracy but also cross-lingual and out-of-domain generalization. For coding problems, we see that MARFT-A-tuned Duo obtains a significant improvement of about 4 scores (~14.75%) and surpasses MARFT-A-tuned Solo by about 2.5 scores (~8.92%). Though we notice that vanilla Trio falls behind the other two, we still see Trio get improved by about 2 scores (~6.81%) by using MARFT-A, which validates the effectiveness of MARFT, regardless of the over-complicated LaMAS design.
Table 4: Evaluation results of Solo, Duo, Trio, and the MARFT-A-tuned ones on CodeForces.
Setup Vanilla MARFT-A
Solo $27.21\pm 2.84$ $28.92\pm 2.07$
Duo $27.45\pm 1.60$ $31.50\pm 1.16$
Trio $26.58\pm 2.47$ $28.39\pm 1.58$
@@ 5.4.1 Optimization Granularity
Figure 6 (Supplementary Figure 8 in Appendix) shows the performance of MARFT-A and MARFT-T while fine-tuning Duo, in mathematical environments. From both subfigures, we see that MARFT-A is overall better than MARFT-T. Specifically, MARFT-A achieves a higher episodic return than MARFT-T in the end and has a more stable and faster improvement. This kind of superiority of MARFT-A over MARFT-T can be theoretically deduced from the inconsistency of their optimality, which is caused by the different MG (Wen et al., 2024b).
@@ 5.5 Future Experiments
Given the preliminary nature of our experiment on the only MATH (Hendrycks et al., 2021), CMATH (Wei et al., 2023) and CodeForces (Penedo et al., 2025), we outline a series of future experiments to further explore the capabilities and robustness of MARFT:
@@ 7 Open Problems
Developing engineering solutions that integrate LLMs and MARL is highly challenging. While there exist user-friendly, efficient, and scalable frameworks such as verl (Sheng et al., 2024), OpenRLHF (Hu et al., 2024) for LLM fine-tuning and MALib (Zhou et al., 2023), ReMA (Wan et al., 2025), PettingLLMs (Zhao et al., 2025b) for MARL training, a comprehensive framework specifically designed for MARFT is notably absent. The creation of such a framework would require significantly more effort than developing either of the aforementioned frameworks individually. This is because it must effectively integrate both LLMs and MARL components, each of which presents its own set of complexities. While this article introduces a MARFT framework, it is still in its early stages and offers significant room for further development and refinement.
@@ 8 Conclusion
In this article, we propose a novel paradigm termed Multi-Agent Reinforcement Fine-Tuning (MARFT), together with a brand-new MG called Flex-MG to better align with the LaMAS optimization in realistic scenarios, and present a universal algorithmic framework, MARFT, that integrates RFT, LaMAS, and MARL to advance the capabilities of LaMAS systems. MARFT addresses several fundamental challenges in adapting conventional MARL to LaMAS, including asynchronous agent interactions, profile-aware architecture design, heterogeneity across agents, and the dynamic organization of multi-agent systems. Through the fusion of RFT with sequential decision-making paradigms, trust-region optimization techniques, and token-level adaptation mechanisms, MARFT enables efficient coordination and policy refinement while preserving the foundational capabilities of pretrained LLMs.
@@ Appendix A Detailed Framework Code Explanation
Abstract MARFT-A training code in the trainer. (For token-level training, we just need to modify the policy loss computation to take token logits as input):
@@
37 critic_grad_norm = $n$.utils.clip_grad_norm_(self.mas.critic.parameters(), self.max_grad_norm)
@@
70 policy_grad_norm = $n$.utils.clip_grad_norm_(self.mas.agents.parameters(), self.max_grad_norm)
@@
76 policy_grad_norm = $n$.utils.clip_grad_norm_(self.mas.agents.parameters(), self.max_grad_norm)
@@ B.2 Hyperparameter Configurations for Experiments
Hyperparameter Value Hyperparameter Value Hyperparameter Value
MAS learning rate 1e-6 critic learning rate 1e-5 use gae True
rollout threads 4 num mini-batch 1 ppo epoch 1
max horizon 1 hidden size 64 episode length 30
optim eps 1e-5 critic loss coef 1 entropy coef 1e-3
gamma 0.99 gae gamma 0.95 max grad norm 0.5
context window 4096 max new tokens 1024 clip param 0.2
Table 7: Hyperparameter configuration for MARFT-T in Figure 6.